johnkt:

Which sequence is modeled by the graph below? (6 points) coordinate plane showing the points 2, 1; 3, 2; 4, 4; and 5, 8

K1NGofPadlet:

This graph doesn't show a specific sequence. It is simply four points plotted on a coordinate plane. The x-coordinates of the points increase from left to right, but I can't infer a specific sequence from these points without additional information, man.

surjithayer:

let \[y=a(b)^x+c\] x=2,y=1 \[1=ab^2+c ...(1)\] x=3,y=2 \[2=ab^3+c ...(2)\] x=4,y=4 \[4=ab^4+c ...(3)\] x=5,y=8 \[8=ab^5+c ...(4)\] (2)-(1),(3)-(2),(4)-(3) gives \[1=ab^2(b-1) ...(5)\] \[2=ab^3(b-1)...(6)\] \[4=ab^4(b-1)...(7)\] divide (6) by (5) b=2 from (5) 1=4a \[a=\frac{ 1 }{ 4 }\] from (1) \[1=\frac{ 1 }{4 }\times 4+c\] c=0 hence \[y=\frac{ 1 }{4 }2^x\]

surjithayer:

or \[y=2^{x-2}\]